OpenStudy (mayataylor):

Medal will be given Which of the following prefixes represents the largest value? A. Giga B. Hector C. Kilo D. Milli

OpenStudy (anonymous):

Giga = 10^9 Hector = 10^2 Kilo = 10^3 Mili = 10^-3

OpenStudy (mayataylor):

Thanks

OpenStudy (mayataylor):

But I found it myself

OpenStudy (anonymous):

Right.